Joanna will graduate from college in two years and she plans to reward herself with a trip to Europe after she graduates. Joanna has saved $12,000 and will make no more deposits to the account during the two years. If Joanna has it in a certificate of deposit earning 4% with quarterly compounding, how much will Joanna have for her trip in two years? Use the calculator: _________________ Use the tables: Table _____ Row ____ Column ____ Factor _____ Answer __________ Use the formula: Answer ____________
Answer #1

Answer:

Given

Deposit D=12000

Annual interest rate =4%

Quarterly interest rate r=4%/4=1%

Number of years =2 years

Number of quarters N=2*4=8 quarters

So Joanna have after two years =D*(1+r)^N=12000*(1+1%)^8=12994.28